OpenStudy (anonymous):

Can Someone Give Me a bit of Help.? I Have a few Questions on Absolute Value! Thanks ^-^ <3

OpenStudy (anonymous):

i can try, whats up?

OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

1) divide both sides by \(-4\)

OpenStudy (anonymous):

If I can i have any help at all id greatly apprciate it

OpenStudy (anonymous):

and i did i got x+5=-4

OpenStudy (anonymous):

2) get rid of the absolute values and solve two separate equations \[x+5=4\\ x+5=-4\]

OpenStudy (anonymous):

divide both sides by -4 then solve for the absolute value. because it is absolute value the answer will be for 4 and -4 of x+5

OpenStudy (anonymous):

o satellite is a faster typer lol

OpenStudy (anonymous):

im confused /.\

OpenStudy (anonymous):

with what part?

OpenStudy (anonymous):

\[-4|x+5|=-16\\ |x+5|=4\] so far so good?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok if you go to solve \(|whatever|=p\) for some positive number \(p\) you do two things, remove the absolute values and get two equations one is \(whatever=p\) and the other is \(whatever=-p\)

OpenStudy (anonymous):

wtf is p?

OpenStudy (anonymous):

in your case \(whatever\) is \(x+5\) so to solve \[|x+5|=4\] you write two equations \[x+5=4\] and also\[x+5=-4\]

OpenStudy (anonymous):

omfg i hate this /.\ fml fml

OpenStudy (anonymous):

in your case \(p=4\)

OpenStudy (anonymous):

p is anything inside the absolute value

OpenStudy (anonymous):

\[|x|=3\] means \(x=3\) or \(x=-3\) \(|x+2|=1\) means \(x+2=1\) or \(x+2=-1\) \(|x+5|=4\) means \(x+5=4\) or \(x+5=-4\)

OpenStudy (anonymous):

non of the answers are 3 -_-

OpenStudy (anonymous):

basically absolute value is asking u for all of the different accountabilities. for ur case that would mean both the positive and negative outcomes

OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

um ur equation ends up being x+5=4 and x+5=-4 so x can be either -1 or -9

OpenStudy (anonymous):

thanks can you help with a few others?

OpenStudy (anonymous):

wait do u understand how i got to that?

OpenStudy (anonymous):

No -.-

OpenStudy (anonymous):

lol well ok u got the |x+5|=4 part right?

OpenStudy (anonymous):

omfg this is stupid af. done fml

OpenStudy (anonymous):

kms -.-

OpenStudy (anonymous):

im getting frusterated.

OpenStudy (anonymous):

hey relax this is not hard. if u understand the concept behind it, it will be better

OpenStudy (anonymous):

sadley too much on my mind to relax....

OpenStudy (anonymous):

ok well take small steps and once u have a full grasp u can apply to other problems

OpenStudy (anonymous):

most math questions are just application of concepts with different numbers

OpenStudy (anonymous):

its just not math i have 101 other things on my mind..

OpenStudy (anonymous):

ok well take a deep breath. most things on ur mind u usually can't control. so I am sure u can narrow that list down=)

OpenStudy (anonymous):

lol if you only knew how long the list actually was....

OpenStudy (anonymous):

ok i understand u have alot on ur mind, but just tackling the small things like concepts can help give u a sense of direction for right now

OpenStudy (anonymous):

idk....thanks anyway